OpenStudy (anonymous):

In the formula , which quantity is measured in units of J/g•°C? A. c B. m C. Q D. DT

OpenStudy (anonymous):

?

OpenStudy (ghazi):

Q=m*c*(T2-T1)...therefore c= Q/m*(T2-T1)....J/g*C

OpenStudy (anonymous):

(A)

OpenStudy (ghazi):

yes :)